OpenStudy (anonymous):

Given f(x) = -2(x + 4)^9 (x + 3)^5 (x-2)^8, find the roots in increasing order. The roots are , , and . To the left of the first root, is the graph of f(x) above or below the x-axis? Answer above or below: . Between the first two roots, is the graph of f(x) above or below the x-axis? Answer above or below: . Between the last two roots, is the graph of f(x) above or below the x-axis? Answer above or below: . After the last root, is the graph of f(x) above or below the x-axis? Answer above or below:

OpenStudy (jack1):

hey @minato ;D so the roots of a function are when y = 0 (aka when does the function cross the x axis) for something that's already factorised like your equation above, it's quite simple to calculate

OpenStudy (anonymous):

Someone who recognized! Wooo! Anyways, go on xP

OpenStudy (jack1):

f(x) = -2(x + 4)^9 (x + 3)^5 (x-2)^8 roots are at f(x) = 0, so: 0 = -2(x + 4)^9 (x + 3)^5 (x-2)^8 now if you have 3 brackets, let's call them A , B , and C f(x) = -2 ( A )^9 ( B )^5 ( C )^8 the possibilities for when f(x) = 0 will occur if: A = 0 as f(x) = -2 ( 0 )^9 ( B )^5 ( C )^8 f(x) = 0 ( B )^5 ( C )^8 and anything times 0 = 0, so it doesn't matter therefore what B and C are. next possibility for f(x) = 0 will occur if: B = 0 as f(x) = -2 ( A )^9 ( 0 )^5 ( C )^8 f(x) = -2 ( A )^9 0 ( C )^8 and anything times 0 = 0, so it doesn't matter therefore what A and C are. final possibility for f(x) = 0 will occur if: C = 0 as f(x) = -2 ( A )^9 ( B )^5 ( 0 )^8 f(x) = -2 ( A )^9 ( B )^5 0 and anything times 0 = 0, so it doesn't matter therefore what A and B are.

OpenStudy (jack1):

so first root: A = 0 and A = (x+4) so 0 = x + 4 -4 = x so f(x) = 0 if x = -4 next root: B = 0 and B = (x+3) so 0 = x + 3 -3 = x @hokage ;) this all makin sense so far dude?

OpenStudy (jack1):

and if so, are you right to solve for final root?

OpenStudy (jack1):

@Yondaime you still with me bro...?

OpenStudy (anonymous):

Sorry about that!

OpenStudy (anonymous):

I kind of get where you were going with this, but the above and below questions still have me at a loss.

OpenStudy (jack1):

that easy enough, there a simple trick to that too, but first, the final root so C = 0 and bracket C = (x - 2) so 0 = (x-2) solve for x dude, then we'll do the above and below stuff, it'll take seconds

OpenStudy (anonymous):

You mean X-2 <-- solve for x? Then x = 2?

OpenStudy (jack1):

perfect! so your roots are x = -4 x = -3 aaaaaaaaaaaand x = 2 

OpenStudy (anonymous):

Alright, cool. I get that part. Thanks for everything so far by the way.

OpenStudy (jack1):

so those points... that's where a graph of your line will = 0 so coordinates are (-4,0) , (-3,0) and (2,0) all good dude hey, naruto fans gotta stick together lol ;D

OpenStudy (anonymous):

Lol, awesome. I like your DP btw ;D DBZ IS THE ROOT OF ALL OF IT.

OpenStudy (jack1):

now: above an below stuff ---> "To the left of the first root, is the graph of f(x) above or below the x-axis? " the first root is (-4, 0 ) so to the left of that, we'll pick a point: how bout x = -5 ? so when x = -5, f(x) = ...? if f(x) = a negative number, it's below the line (x axis) if f(x) = a positive number, it's above the line if f(x) = 0, it's still on the line so i'm gonna use an equation solver so we dont have to do the heavy maths here, but what you do is sub x = -5 into the big equation to get your answer for f(x) SO: f(x) = -2(x + 4)^9 (x + 3)^5 (x-2)^8 at x = -5 f(-5) = -2(-5 + 4)^9 (-5 + 3)^5 (-5-2)^8 = -2(-1)^9 (-2)^5 (-7)^8 = -368947264 which is a freaking huge NEGATIVE number, so f(x) is below the line

OpenStudy (jack1):

next one: "Between the first two roots, is the graph of f(x) above or below the x-axis?" so pick any number between x = -4 and x = -3 (lets go x = -3.5) plug it into the equation as per the previous answer: f(x) = -2(x + 4)^9 (x + 3)^5 (x-2)^8 at x = -3.5 f(-3.5) = -2(-3.5 + 4)^9 (-3.5 + 3)^5 (-3.5-2)^8 answer: f(-3.5) = 102.214 it's a positive number, so f(x) is ABOVE the line (graph of the function attached)

OpenStudy (jack1):

next: Between the last two roots, is the graph of f(x) above or below the x-axis? so pick any number between x = -3 and x = -2 (lets go x = 0) plug it into the equation as per the previous answer: (seeing a pattern here ;D ) f(x) = -2(x + 4)^9 (x + 3)^5 (x-2)^8 at x = 0 f(0) = -2(0 + 4)^9 (0 + 3)^5 (0-2)^8 f(0) = -2( 4)^9 ( 3)^5 (-2)^8 f(0) = -32614907904 massive number, but more importantly: NEGATIVE, so f(0) is BELOW the line

OpenStudy (jack1):

cool dude, now u go nuts on the last question: "After the last root, is the graph of f(x) above or below the x-axis?"

OpenStudy (jack1):

so pick any number higher than x = 2

OpenStudy (anonymous):

I see..wow, thanks a lot man. You took your time and really got it to me. Thank you so much.

OpenStudy (jack1):

all good dude, always welcome hey ;D

OpenStudy (jack1):

now you can either look at the graph or do the maths for the last Q, up to u slaters ;D